OpenStudy (celinegirl):

Choose the correct slope of the line that passes through the points (−4, 8) and (−3, −6). 7 −14 0 14

OpenStudy (abdullahm):

\(\sf\large Slope = \frac{y_2-y_1}{x_2-x_1} = \frac{8-(-6)}{-4-(-3)}=?\)

OpenStudy (mrnood):



OpenStudy (abdullahm):

for the points \(\sf\large (x_1,y_1)\) and \(\sf\large (x_2,y_2)\)

OpenStudy (mrnood):

@celinegirl so - you have all the hints and clues you need what is the answer?

OpenStudy (celinegirl):

Is it 7?

OpenStudy (abdullahm):

nope

OpenStudy (celinegirl):

Hm.. is it 14? @AbdullahM

OpenStudy (abdullahm):

Almost correct, it's actually -14

OpenStudy (abdullahm):

\(\sf\large Slope = \frac{y_2-y_1}{x_2-x_1} = \frac{8-(-6)}{-4-(-3)}=\frac{8+6}{-4+3} = \frac{14}{-1} = -14\)

OpenStudy (abdullahm):

@celinegirl

OpenStudy (celinegirl):

@AbdullahM Oh!! okay that makes sense. Thanks lol(:

OpenStudy (abdullahm):

np :)